# GUICourse: From General Vision Language Model to Versatile GUI Agent

### v1

@@ Abstract
Utilizing Graphic User Interface (GUI) for human-computer interaction is essential for accessing a wide range of digital tools. Recent advancements in Vision Language Models (VLMs) highlight the compelling potential to develop versatile agents to help humans finish GUI navigation tasks. However, current VLMs are challenged in terms of fundamental abilities (OCR and grounding) and GUI knowledge (the functions and control methods of GUI elements), preventing them from becoming practical GUI agents. To solve these challenges, we contribute GUICourse, a suite of datasets to train visual-based GUI agents from general VLMs. First, we introduce the GUIEnv dataset to strengthen the OCR and grounding capabilities of VLMs. Then, we introduce the GUIAct and GUIChat datasets to enrich their knowledge of GUI components and interactions. Experiments demonstrate that our GUI agents have better performance on common GUI tasks than their baseline VLMs. Even the small-size GUI agent (with 3.1B parameters) can still work well on single-step and multi-step GUI tasks. Finally, we analyze the different varieties in the training stage of this agent by ablation study. Our source codes and datasets are released at https://github.com/yiye3/GUICourse.
@@ 1 Introduction
Graphical User Interfaces (GUIs) serve as a pivotal medium for facilitating human-computer interactions, playing a pivotal role across diverse applications. Different GUI systems (e.g. websites and smartphones) usually use consistent icons and operation logic, which makes users ignore the complexity of the underlying processes and promotes skill transfer between different systems. GUI agents, which are constructed for GUI navigation tasks, not only liberate humans from tedious and repetitive operations, but can also be regarded as an important method for using GUI tools, thereby expanding the capabilities of large language models (LLMs).
Recent works have explored training GUI agents in various scenarios, such as computer control [14], web browsing [7, 44], and smartphone navigation [31, 36, 39, 43]. These agents can be classified based on their input modalities into textual-based (e.g., HTML, source code) GUI agents and visual-based (e.g., screenshots) GUI agents. We argue that visual-based GUI agents have two advantages over textual-based methods: (1) Ease of acquisition: Obtaining screenshots of GUI systems is straightforward, whereas accessing structured text, such as the code behind the iOS system, is challenging. (2) Improved transferability: GUIs typically utilize similar visual elements to denote similar functions across different systems.
The performance of vision-based agents heavily relies on the fundamental capabilities and internal knowledge of current Vision Language Models (VLMs). Although VLMs [20, 26, 52, 17, 13] demonstrate impressive capabilities in various tasks such as image captioning [18] and visual question answering [21], these capabilities may not suffice for developing efficient visual-based GUI agents. For example, we investigate Qwen-VL-Chat, a powerful open-sourced VLM, on its ability to finalize GUI navigation instructions. As illustrated in Figure 1, it shows the two challenges of current VLMs: (1) Their OCR and grounding abilities are not supported to locate the website elements or designated text accurately; (2) They lack a comprehensive understanding of the functions and control mechanisms of website widgets.
@@
In this work, We contribute GUICourse, a group of complete datasets to help improve VLMs’ fundamental abilities and GUI knowledge. GUICourse is composed of three datasets: (1) GUIEnv, a large-scale dataset for improving VLMs’ OCR and grounding abilities, including 10M website page-annotation pairs as pre-training data and 0.7M region-text QA pairs as SFT data; (2) GUIAct, a GUI navigation dataset in website and android scenarios for enhancing VLMs’ knowledge of GUI systems, including 67k single-step and 15k multi-step action instructions. (3) GUIChat, a conversational dataset for improving the interaction skills of GUI agents, including 44k single-turn QA pairs and 6k multi-turn dialogues with text-rich images and bounding boxes. Based on the proposed dataset, we train several GUI agents based on different VLMs (Qwen-VL, Fuyu-8B, and MiniCPM-V). We demonstrate that our GUI agents have better performance on Mind2Web and AITW than their baseline VLMs. Then, we conduct ablation studies on our GUI agent MiniCPM-GUI, comparing the impact of different GUIEnv pre-training data amounts, image resolutions, and mixed GUIChat data. We demonstrate that GUIEnv data is effective in improving VLMs’ OCR and grounding abilities. These abilities, as well as high resolution, significantly increase the performance of GUI tasks.
@@ 2.1 Vision Language Models
Currently, most Vision-Language Models (VLMs) use a bridge module to connect a vision encoder, such as CLIP-ViT [8, 28], and a language model, such as Vicuna [38] and Qwen [3]. The bridge module maps the visual features from the vision encoder into the embedding space of the large language model with an MLP layer, exemplified by LLaVA [21], or an attention-based resampler, such as BLIP-2 [18] and Qwen-VL [4]. Uniquely, Fuyu-8B [5] removes the external vision encoder and uses a decoder-only transformer architecture with a unified space for textual tokens and pixel information, enabling Fuyu-8B to process images of any resolution. Based on this, subsequent research efforts try to enhance VLMs in various domains, including visual detail recognition [40, 49], reliable responses [37, 47], object grounding [26], multilingual capabilities [13] and model efficiency [53]. However, many VLMs still have shortcomings when (1) recognize small-size text in high-resolution images, and (2) provide the accurate location of the designated area at the pixel level. In this work, we provide the GUIEnv dataset to improve the VLMs’ OCR and grounding abilities with high-resolution website screenshots.
@@ 2.2 GUI Agents
GUI agents are expected to help humans finish tasks on different GUI systems, such as simplified websites [33, 19], simulated environments [44, 51], or real-world websites and smartphones [7, 31]. Nowadays, the GUI agents are developing quickly due to the Large Models (LMs), Big Data, and data-driven methods. Some agents [10, 34, 23, 15, 41, 50, 43] using the train-free method and depending on closed-sourced LMs (e.g., GPT4 and GPT-4V [25]) to achieve GUI navigation with strategy design and prompt engineering. Also, some agents [14, 24, 27] finish GUI tasks based on small-size open-sourced models, using data-driven methods (e.g., supervised fine-tuning).
@@
Perception and control are the two important points for GUI agents. According to the perception modalities of environments, the GUI agents can split into text-only [11, 7], image-text combined [14, 9], and vision-only [32, 48, 12] methods. As for the control, GUI agents can split into position-free and position-needed methods. The position-free methods [7, 44] generally provide indexes for every element in the GUI systems, and agents can finish actions by multiple choices. While the position-needed methods [32, 48, 12] mean agents need to give the position information to finish some actions, such as "click" and "tap". In this work, we aim to train GUI agents with vision-only inputs and position-needed actions using data-driven methods.
To train a visual-based GUI agent using data-driven methods, lots of GUI navigation data with visual modal information (e.g., screenshots) are needed. Some works provide GUI navigation datasets including visual modal information, but their datasets usually have one or more of the following problems: (1) The GUI environments are too easy, and far from real-world scenarios [19, 14]. (2) The datasets are focused on a narrow domain or scenario [44, 35, 31]. (3) The size of datasets is too small to train GUI agents [7]. In this work, we contribute the GUIAct dataset to solve these three problems. Besides that, we also provide the GUIChat dataset to improve the GUI agents’ natural language interaction abilities.
3 GUICourse Construction
@@
In this section, we introduce GUICourse, a group of datasets for improving VLMs’ OCR and grounding abilities, enhancing VLMs’ GUI system knowledge, and helping GUI agents’ interaction. As shown in Table 1, our GUICourse is the first large-scale corpus for the entire pipeline to train GUI agents from general VLMs, which includes GUIEnv, GUIAct, and GUIChat.
Table 1: Statistics of the GUICourse compared with existing datasets. The columns indicate: whether data includes generating and understanding the content in designated regions and full pages (Region and Page), the number of training samples (Num.), whether real-world scenarios are used (Real.), the number of instructions in website and android scenarios (Web Inst. and And. Inst.), whether there are text-rich images and bounding boxes in dialogues (Text-Rich and w/ Bbox).
Dataset OCR & Grounding GUI Knowledge Interaction
Region Page Num. Real. Web Inst. And. Inst. Text-Rich w/ Bbox
MiniWoB++ [19] - 100 -
WebShop [44] - 12k -
META-GUI [36] - - 1,125
MIND2WEB [7] - 2,350 -
AITW [31] - - 715k
WEBLINX [22] - 2,337 -
SeeClick [6] <1M - - -
ScreenAI [2] - - - -
Ferret-UI [46] <1M - - -
GUICourse (Ours) 10.7M 73k 9,157
3.1 GUIEnv: A Large Scale Dataset for Fundamental Abilities
@@
OCR and grounding are two fundamental abilities for visual-based GUI agents’ perception and control. To augment these abilities, we create a large-scale dataset called GUIEnv with about 10M pre-training data GUIEnv-global and 0.7M SFT data GUIEnv-local. As shown in Figure 2, each sample in GUIEnv-global is a long text with all the describable content on the full page, including text, grounding information, and layout sequence. Each sample in GUIEnv-local is a QA pair on a designed region as the "text2bbox" or "bbox2text" task.
@@
We collected 4M URLs from the C4 [29] and employed Playwright [^1] https://github.com/microsoft/playwright for rendering, ultimately producing 10M annotated screenshots[16]. The 10M annotated screenshots correspond to the 10M samples of the GUIEnv-global data. As for the GUIEnv-local data, we select about 50k website screenshots with their elements from C4. These screenshots are usually huge due to they include the complete content of the websites. We process the data in three steps: (1) crop them to smaller partitions less than $1920\times 1080$ resolution. (2) remove images with elements less than 10. (3) randomly simply 10 elements with texts and positions in each screenshot as the data of "text2bbox" and "bbox2text" tasks. Finally, we acquired 0.7M instructions for the GUIEnv-local data.
3.2 GUIAct: A Multi-Scenario Dataset for GUI Knowledge
GUI systems’ knowledge, which includes the functions and the control methods of GUI elements, is important for GUI agents. Aiming at bridging the gap between VLMs and GUI agents, we contribute the GUIAct dataset to enhance GUI knowledge. The GUIAct dataset includes GUI navigation tasks in two scenarios (website and smartphone), which can be split into three partitions: "web-single", "web-multi", and "smartphone". GUI navigation tasks require agents to interact with GUI environments using actions, according to a task instruction, as shown in Figure 3. In this work, we only consider the GUI agents to interact with image-only environments. We only use the screenshots of GUI systems in the viewport, without any unseen elements, which use the same environments as humans do. Different GUI systems generally have diverse actions (e.g. "click" is common on websites while "tap" is common on smartphones). We define a unified action space for different GUI systems, including 11 types of actions, and the details are shown in Appendix A.1.
@@ 3.2 GUIAct: A Multi-Scenario Dataset for GUI Knowledge
We contribute the GUIAct (web-single) dataset in four steps: (1) Websites selection. We use GPT-4 to gather diverse scenarios (e.g., online shopping and E-learning) and URLs of using the internet. Then, we use these URLs as seeds to expand new websites by their hyperlinks. We acquire 50 domains and 13k websites in this step. (2) Captures acquisition. We use web snapshot tools to respond to website HTML, interactive elements, and screenshots based on their URLs. (3) LLM-Auto Annotation. We use GPT-4V to obtain single-step instruction-action pairs for each website. We give two images to GPT-4V in each request: an origin screenshot and a revised screenshot with interactive element identification. We get about 70k instruction-action pairs based on the 13k website screenshots. (4) Data checking. We hire annotators to check the automatic instruction-action pairs’ quality. The annotators are requested to revise the inaccurate items or abandon them if they are hard to revise. After that, the accuracy of GUIAct (web-single) data improves from 55% to 92%, based on our sampled results. Finally, we get around 67k single-step action instructions in website scenarios, corresponding to 67k training samples.
We contribute the GUIAct (web-multi) dataset in three steps: (1) Websites selection. We concept 8 top-level web scenarios for information searching: Shopping, Diet, Stay, Travel, Education, Health, Job, and Entertainment. Based on the 8 top-level scenarios, we collect 32 sub-scenarios and 121 well-known websites. (2) Questions acquisition. We use LLMs (GPT-3.5 and Claude2) to generate instructions for each website. We collect 8k high-level instructions for searching special information on the corresponding websites. To finish each instruction needs multi-turn interaction with the website. (3) Crowd-sourced Annotation. We develop an annotation tool as a web browser plugin. Then, we hire annotators to execute operators to finish the instructions on the corresponding websites. If the instructions are inaccurate or have no answer, the annotators will revise or abandon them. Finally, We get 5,696 multi-step action instructions in website scenarios, corresponding to 44k training samples.
We contribute the GUIAct (smartphone) dataset using a subset of the AITW [31] dataset. We select the data with the "General" tag and filter out the screenshots without the bottom navigation bar (because we will convert the "go back" and "go home" actions to "tap"). Then, we convert the origin actions to our action space. Finally, we get 9,157 multi-step action instructions in smartphone scenarios, corresponding to 67k training samples.
3.3 GUIChat: A Text-Rich Multi-Modal Dataset for Interaction

### v2

@@ Abstract
Utilizing Graphic User Interfaces (GUIs) for human-computer interaction is essential for accessing various digital tools. Recent advancements in Vision Language Models (VLMs) reveal significant potential for developing versatile agents that assist humans in navigating GUIs. However, current VLMs face challenges related to fundamental abilities, such as OCR and grounding, as well as a lack of knowledge about GUI elements functionalities and control methods. These limitations hinder their effectiveness as practical GUI agents. To address these challenges, we introduce GUICourse, a series of datasets for training visual-based GUI agents using general VLMs. First, we enhance the OCR and grounding capabilities of VLMs using the GUIEnv dataset. Next, we enrich the GUI knowledge of VLMs using the GUIAct and GUIChat datasets. Our experiments demonstrate that even a small-sized GUI agent (with 3.1 billion parameters) performs effectively on both single-step and multi-step GUI tasks. We further fine-tune our GUI agents on other GUI tasks with different action spaces (AITW and Mind2Web), and the results show that our agents are better than their baseline VLMs. Additionally, we analyze the impact of OCR and grounding capabilities through an ablation study, revealing a positive correlation with GUI navigation ability. Our source codes and datasets are released at https://github.com/RUCBM/GUICourse.
GUICourse: From General Vision Language Model to Versatile GUI Agent
@@ 1 Introduction
Graphical User Interfaces (GUIs) are a pivotal medium for facilitating human-computer interactions, playing a crucial role across diverse applications. GUI agents are designed to complete complex tasks in these GUI systems, which can liberate humans from tedious and repetitive operations. Different GUI systems (e.g., websites and smartphones) usually use icons and images to convey specific information, making them suitable to be processed by visual-based agents. On the other hand, obtaining screenshots of GUI systems is straightforward, whereas accessing structured text is challenging (such as the code behind the iOS system). In this work, we focus on training visual-based GUI agents from general VLMs.
The performance of vision-based agents heavily relies on the fundamental capabilities and internal knowledge of their baseline VLMs. Although VLMs (Liu et al., 2023a; Peng et al., 2023; Zhu et al., 2023; Li et al., 2023a; Hu et al., 2023) demonstrate impressive capabilities in various tasks such as image captioning (Li et al., 2023b) and visual question answering (Liu et al., 2023b), these capabilities may not suffice for developing efficient visual-based GUI agents. For instance, we investigate the Qwen-VL-Chat, a powerful open-sourced VLM, on its ability to finalize GUI navigation instructions. As illustrated in Figure 1, it shows the two challenges of current VLMs: (1) Their OCR and grounding abilities are not supported to locate the website elements or designated text accurately; (2) They lack a comprehensive understanding of the functions and control mechanisms of website widgets. To solve these problems, we provide a pipeline for training GUI agents based on general VLMs. This involves enhancing their OCR and grounding capabilities, followed by training them to perform GUI navigation tasks and engage in related conversations.
@@
Our approach mainly employs data-driven methods for training diverse agents. To cultivate versatile GUI agents, we require a substantial amount of high-quality GUI navigation data enriched with visual information (e.g., screenshots). While previous works have contributed various datasets, they often suffer from one or more shortcomings: (1) the GUI environments may be overly simplistic and disconnected from real-world scenarios Liu et al. (2018); Humphreys et al. (2022); (2) the datasets tend to focus on narrow domains or scenarios Yao et al. (2022); Sun et al. (2022a); Rawles et al. (2023); and (3) the overall size of the datasets may be insufficient for effective training of GUI agents Deng et al. (2023). In this work, we primarily introduce the GUIAct dataset, which addresses these challenges by providing 67k human-verified single-step and 5,696 human-annotated multi-step GUI navigation instructions across various websites. Additionally, we present the GUIEnv dataset, comprising 10 million website page-annotation pairs, and the GUIChat dataset, featuring 44k single-turn QA pairs and 6k multi-turn dialogues alongside website screenshots, to support our training pipeline.
In summary, our contributions are as follows:
(1) We establish a comprehensive pipeline and a series of datasets for building GUI agents based on VLMs. This pipeline consists of two learning stages aimed at enhancing fundamental abilities and acquiring knowledge specific to GUI systems.
(2) We train several visual-based GUI agents according to the pipeline and datasets. We find that even small-sized agents perform effectively in GUI navigation tasks. Additionally, we demonstrate that integrating OCR and grounding data during the pre-training phase is helpful for these tasks.
@@ 2.1 Vision Language Models
Currently, most Vision-Language Models (VLMs) use a bridge module to connect a vision encoder, such as CLIP-ViT (Dosovitskiy et al., 2020; Radford et al., 2021), and a language model, such as Vicuna (Touvron et al., 2023) and Qwen (Bai et al., 2023a). The bridge module maps the visual features from the vision encoder into the embedding space of the large language model with an MLP layer, exemplified by LLaVA (Liu et al., 2023b), or an attention-based resampler, such as BLIP-2 (Li et al., 2023b) and Qwen-VL (Bai et al., 2023b). Uniquely, Fuyu-8B (Bavishi et al., 2023) removes the external vision encoder and uses a decoder-only transformer architecture with a unified space for textual tokens and pixel information, enabling Fuyu-8B to process images of any resolution. Based on this, subsequent research efforts try to enhance VLMs in various domains, including visual detail recognition (Wu and Xie, 2023; Zhang et al., 2024), reliable responses (Sun et al., 2023b; Yu et al., 2023), object grounding (Peng et al., 2023), multilingual capabilities (Hu et al., 2023) and model efficiency (Zhu et al., 2024). The current VLMs normally have OCR ability when recognizing sparse characters in small-size images (Mishra et al., 2019; Singh et al., 2019) or dense text with similar fonts (Blecher et al., 2023). However, they still have shortcomings when recognizing texts with different fonts in large screenshots and providing their accurate locations at the pixel level. In this work, we provide the GUIEnv dataset to improve the VLMs’ OCR and grounding abilities with high-resolution website screenshots.
@@ 2.2 GUI Agents
GUI agents are expected to help humans finish tasks on different GUI systems, such as simplified websites (Shi et al., 2017; Liu et al., 2018), simulated environments (Yao et al., 2022; Zhou et al., 2023), or real-world websites and smartphones (Deng et al., 2023; Rawles et al., 2023). Nowadays, the GUI agents are developing quickly due to the Large Models, Big Data, and data-driven methods. Some agents (Gao et al., 2023; Sun et al., 2023a; Ma et al., 2023; Kim et al., 2023; Xu et al., 2021; Zheng et al., 2023; Yang et al., 2023) using the train-free method and depending on closed-sourced Large Models, such as GPT4 and GPT-4V (OpenAI, 2023), to achieve GUI navigation with strategy design and prompt engineering. Also, some agents (Humphreys et al., 2022; Nakano et al., 2021; Qin et al., 2023) finish GUI tasks based on small-size open-sourced models, using data-driven methods (e.g., SFT).
@@
Perception and control are the two important points for GUI agents. According to the perception modalities of environments, the GUI agents can split into text-only (Gur et al., 2023; Deng et al., 2023), image-text combined (Humphreys et al., 2022; Furuta et al., 2023), and vision-only (Shaw et al., 2023; Zhan and Zhang, 2023; Hong et al., 2023) methods. As for the control, GUI agents can split into position-free and position-needed methods. The position-free methods Deng et al. (2023); Yao et al. (2022) generally provide indexes for every element in the GUI systems, and agents can finish actions by multiple choices. While the position-needed methods Shaw et al. (2023); Zhan and Zhang (2023); Hong et al. (2023) mean agents need to give the position information to finish some actions, such as "click" and "tap". Using text-based perception or position-free actions requires extracting element information from GUI systems. That might be difficult to obtain due to system permission restrictions. In contrast, vision-based information (e.g., screenshots) is readily available, and position-based actions align with the original design of GUI interactions. In this work, we aim to train GUI agents with vision-only inputs and position-based actions.
3 GUICourse
@@
In this section, we introduce GUICourse, a series of datasets for improving VLMs’ OCR and grounding abilities, enhancing VLMs’ GUI knowledge, and helping GUI agents’ interaction. As shown in Table 1, our GUICourse includes GUIEnv, GUIAct, and GUIChat three partitions. Examples from the GUICourse datasets are provided in Appendix C. We employ a combination of LLM-based auto-annotation and human annotation to construct our datasets. More details of the annotation process (e.g., the prompts and costs) are shown in Appendix A.2 and Appendix A.3.
Name Subset Source Num of Ins.
GUIEnv global C4 10M
local C4 0.7M
GUIAct web-single GPT-4V/Human 67k
web-multi Human 5,696
smartphone AITW-general 9,157
GUIChat single-turn GPT-4 44k
multi-turn GPT-4 6k
Table 1: The overview of datasets we contributed for multi-stage GUI Agent training.
Dataset Dom. Env. Inst. Avg. Turns Inst. Level Scenarios
MiniWoB++ (Liu et al., 2018) - 100 100 3.6 Low-level Simplified Mob./Webs
Rico (Deka et al., 2017) 27 9,772 10,811 - - (Traces) Mob. Apps
UIBert (Bai et al., 2021) - - 16,660 1.0 Low-level Mob. Apps
UGIF (Venkatesh et al., 2022) - 12 523 5.3 High & low Mob. Apps
WebShop (Yao et al., 2022) 1 1 12k 11.3 High-level Shopping Webs
RUSS (Xu et al., 2021) - 22 80 - High & low Real-world Webs
PixelHelp (Li et al., 2020) 4 4 187 - High & low Mobile Apps
META-GUI (Sun et al., 2022b) 6 11 1,125 4.3 High-level (Dialogues) Mob. Apps
MoTIF (Burns et al., 2022) 15 125 756 - High & low Mobile Apps
MIND2WEB (Deng et al., 2023) 5/31 137 2,350 7.3 High-level Real-world Webs
AITW (Rawles et al., 2023) - 357 715k 6.5 High & low Android/Apps
WEBLINX (Lù et al., 2024) - 155 2,337 43.0 High-level (Dialogues) Webs
AndroidControl (Li et al., 2024) - 833 15,283 4.8 High & low Mobile Apps
GUI-World (Chen et al., 2024) - - 12,379 - High-level Web./Mob./Desk./XR
GUIAct (web-single) 50 13k 67k 1.0 Low-level Real-world Webs
GUIAct (web-multi) 8/32 121 5,696 7.9 High-level Real-world Webs
Table 2: Statistics of the GUIAct dataset compared with existing datasets. The columns indicate: the number of app/website domains, environments, and instructions (Dom., Env., and Inst.), the average number of turns per instance (Avg. Turns), instruction level, and scenarios. Our main contributions are the two largest subsets in website scenarios with both single-step and multi-step instructions.
3.1 GUIEnv: A Large Scale Dataset for OCR and Grounding Abilities
@@
OCR and grounding are two fundamental abilities for visual-based GUI agents’ perception and control. To augment these abilities, we create a large-scale dataset called GUIEnv, which includes two subsets named GUIEnv-global and GUIEnv-local. GUIEnv-global has 10M samples for the pre-training, and each sample is a long text with all the describable content on the full page, including text, grounding information, and layout sequence. GUIEnv-local has 0.7M samples for the SFT, and each sample is a QA pair on a designated region as the "text2bbox" or "bbox2text" task. The samples of the two subsets are shown in Figure 14.
@@
We collected 4M URLs from the Cleaned Common Crawl Corpus (Raffel et al., 2019) which is referred as C4. Then, we employed Playwright  [^1] https://github.com/microsoft/playwright for rendering, ultimately producing 10M annotated screenshots (Lee et al., 2023). These screenshots correspond to the 10M samples of the GUIEnv-global data. As for the GUIEnv-local data, we select about 50k website screenshots with their elements from C4. These screenshots are usually huge because they include the websites’ complete content. We process the data in three steps: (1) crop them to smaller partitions less than $1920\times 1080$ resolution, (2) remove images with elements less than 10, (3) randomly sample 10 elements with texts and positions in each screenshot as the data of "text2bbox" and "bbox2text" tasks. Finally, we acquired 0.7M instructions for the GUIEnv-local data.
3.2 GUIAct: A GUI Navigation Dataset with Single-step and Multi-step Instructions
We argue that GUI agents can learn GUI knowledge, such as the functions and the control methods of GUI elements from GUI navigation tasks. We provide the GUIAct dataset, which can be split into three partitions: single-step web navigation, multi-step web navigation, and smartphone navigation. Our main contributions are two subsets in website scenarios: a single-step website navigation subset with 67k instructions and a multi-step website navigation subset with 5,696 instructions. As shown in Table 2, although there are many similar navigation tasks in mobile application scenarios, the two subsets are the largest navigation datasets in real-world website scenarios with both single-step and multi-step instructions. An example of our multi-step instruction is shown in Figure 2. The GUI agents only receive screenshots of GUI systems in the viewport, without any unseen elements, and output the action with position. We define a unified action space for different GUI systems, including 11 types of actions, and the details are shown in Appendix A.1. To enhance the scenario diversity of our dataset, we also convert a part of the samples in AITW depending on our action space and action format. We use the abbreviations "web-single", "web-multi", and "smartphone" to denote the three subsets in the subsequent context.
@@ 3.2 GUIAct: A GUI Navigation Dataset with Single-step and Multi-step Instructions
We contribute the GUIAct (web-single) dataset in four steps: (1) Websites selection. We use GPT-4 to gather 50 different scenarios (e.g., online shopping and E-learning) and hundreds of URLs. Then, we use these URLs as seeds to expand new websites by their hyperlinks, acquiring about 13k websites. (2) Captures acquisition. We use web snapshot tools to respond to website HTML, interactive elements, and screenshots based on their URLs. (3) LLM-Auto Annotation. We use GPT-4V to obtain single-step instruction-action pairs for each website. We give two images to GPT-4V in each request: an origin screenshot and a revised screenshot with interactive element identification. We get about 70k instruction-action pairs based on the 13k website screenshots. (4) Data checking by human. We hire annotators to check the automatic instruction-action pairs’ quality. The annotators are requested to revise the inaccurate items or abandon them if they are hard to revise. After that, the accuracy of GUIAct (web-single) data improves from 55% to 92%, based on our sampled results. Finally, we get around 67k single-step action instructions in website scenarios, corresponding to 67k training samples.
We contribute the GUIAct (web-multi) dataset in three steps: (1) Websites selection. We concept 8 top-level common web scenarios: Shopping, Diet, Stay, Travel, Education, Health, Job, and Entertainment. Based on these scenarios, we collect 32 sub-scenarios and 121 well-known websites. (2) Questions acquisition. We collect 8k high-level instructions related to searching for specific information on the corresponding websites. (3) Crowd-sourced Annotation. We develop an annotation tool as a web browser plugin. Then, we hire annotators to execute operators to finish the instructions on the corresponding websites. If the instructions are inaccurate or have no answer, the annotators will revise or abandon them. Finally, we get 5,696 multi-step action instructions in website scenarios, corresponding to 44k training samples.
We contribute the GUIAct (smartphone) dataset using a subset of the AITW (Rawles et al., 2023) dataset. We select the data with the "General" tag and filter out the screenshots without the bottom navigation bar (we will convert the "go back" and "go home" actions to "tap"). Then, we convert the origin actions to our action space. Finally, we get 9,157 multi-step action instructions in smartphone scenarios, corresponding to 67k training samples.
3.3 GUIChat: A Text-Rich Multi-Modal QA Dataset for Interaction